# Spins of Black Holes in X-ray Binaries and the Tension with the Gravitational Wave Measurements

### v2

@@ 3.1.2 The continuum method: standard models
Notes: A version of bhspec covering $a_{*}$ of 1-1–0.99 is available from Shane Davis upon request. The normalization should be kept at unity in all of the models except in bhspec, where it is tied to the source distance. The $f_{\rm col}$ of kerrbb2 is obtained by fitting its values in kerrbb to the corresponding spectra generated from bhspec.
We stress that an important issue regarding the continuum method concerns the disk’s local rest-frame spectra. Table 1 lists the method used by the main available model and the ranges of the main parameters, $a_{*}$, the viscosity parameter, $\alpha$, and the Eddington ratio, $L_{\rm d}/L_{\rm Edd}$, where $L_{\rm d}$ is the disk luminosity, and $L_{\rm Edd}$ is the Eddington luminosity for hydrogen. A commonly used model kerrbb (Li et al., 2005) approximates the local spectra as blackbodies with a color correction, $f_{\rm col}$ (Shimura and Takahara, 1995; Davis and El-Abd, 2019), where this factor is defined as the ratio of the color temperature to the effective one, $f_{\rm col}\equiv T_{\rm color}/T_{\rm eff}$. The value of $f_{\rm col}$ for a given disk is uncertain, though usually in the range of 1.5–2 (Davis and El-Abd, 2019). The same parametrization is used by the model kynbb [^1] https://projects.asu.cas.cz/stronggravity/kyn (Dovčiak et al., 2004b), which also allows the inner disk radius to be different from the ISCO radius, $R_{\rm ISCO}$.
@@
Finally, we notice that the simple nonrelativistic model of the disk continuum, diskbb (Mitsuda et al., 1984), is often used to roughly estimate the spin. The model gives an approximate (and proportional to the distance) estimate of the disk inner radius, which can be converted to the spin if the mass and distance are known. This model follows the formulation of Shakura and Sunyaev (1973) except that it does not include the zero-stress boundary condition at the ISCO, $(1-\sqrt{R_{\rm ISCO}/R})$. The distance, $D$, inclination, $i$, and the inner radius are related to the diskbb normalization, $N_{\rm dbb}$, by
@@ 3.1.3 The continuum method: problems and alternatives
An example of this class of models is the so-called magnetically elevated disks (Begelman and Pringle, 2007; Begelman and Silk, 2017; Mishra et al., 2020). Importantly, they have lower gas densities at the effective photosphere, which enhances electron scattering and could lead to a harder spectrum than from a disk without magnetic pressure support. For this reason, their average color correction could be significantly larger than the standard range of $f_{\rm col}\approx 1.5$22. Unfortunately, to our knowledge, there have been no radiative transfer calculations of the spectra from magnetically elevated disks as of yet. Still, there have been recent advancements in numerical simulations of accretion flows, including more realistic modeling of radiation transfer (e.g., Sądowski and Narayan, 2016; Huang et al., 2023; Fragile et al., 2023) as well as more sophisticated tools for generating spectral predictions (e.g., Kinch et al., 2019; Mills et al., 2024; Roth et al., 2025).
Alternatively, a significant part of the total pressure support could be provided by a radial magnetic field, in addition to the toroidal one (Lančová et al., 2019), resulting in a thicker disk, with a photosphere that is hotter than that of a standard disk, see Figure 2 in Wielgus et al. 2022). General relativistic radiative magnetohydronamic (GRRMHD) simulations of these so-called puffy disk models have been considered in the regime $\dot{M}_{\rm accr}\approx\dot{M}_{\rm Edd}$ (where we define $\dot{M}_{\rm Edd}\equiv L_{\rm Edd}/0.1c^{2}$). An example of the vertical structure of a puffy disk is shown in Figure 4, where the photosphere is located at around $z\approx r$ (solid white line at $\tau_{\rm T}=1$), while the disk scale height is $z\approx 0.1\,r$ (dashed white line). The presence of a hotter and geometrically thick photosphere reaching to the BH event horizon results in a significant overestimate of the BH spin. When the kerrbb and slimbh models were used on synthetic spectra obtained from the simulations assuming $a_{*}=0$, the fitted spin values ranged from $a_{*}\approx 0.5$ to $\approx$0.9 (Wielgus et al., 2022; Lančová et al., 2023). While these numerical simulations, like all others, come with numerous caveats, this case study is illuminating because it highlights how the more complex physical behavior seen in simulations (and potentially occurring in nature) can lead to significant biases in BH spin estimates, even when interpreted with state-of-the-art spectral models.
@@
The similarity of the disk and coronal electron temperatures also prevents the formation of a power law spectral component due to Comptonization of the disk photons by those electrons. Instead, the overall spectral shape is modified. Together with the Galactic absorption being strong at $\lesssim$ 1 keV, these effects hinder placing effective spectral constraints on the existence of soft X-ray excesses in XRBs. Nevertheless, soft X-ray excesses have been found in some BH X-ray binaries (BH XRBs), particularly in the soft state of GRS 1915+105 (a BH X-ray binary with a low-mass donor, hereafter LMXB), which was fitted by a hybrid Comptonization model including a low-temperature, high optical depth thermal component (Zdziarski et al., 2001). A similar result was found for the very high state of GRO J1655–40 (Kubota et al., 2001).
Following those results, models with a warm coronae were fitted by Belczyński et al. (2024) and Zdziarski et al. (2024a, b, 2025) to the soft states of LMC X-1, Cyg X-1, M33 X-7 and GX 339–4. As we discuss below, this model provides statistically good fits and significantly alters the fitted values of the spin. However, the one-zone Comptonization model used so far for the warm corona (Zdziarski et al., 2020) does not account for the detailed corona structure presented in Gronkiewicz and Różańska (2020).
@@
Then, both analytical work and numerical simulations show that the viscous $\alpha$ parameter cannot be taken as constant, typically rising steeply near the ISCO and reaching a sharp maximum within the plunging region (Krolik, 1999; Agol and Krolik, 2000; Sorathia et al., 2012; Penna et al., 2013; Abramowicz et al., 1996; Jiang et al., 2019). Thus, this increase of stress will also affect the flow within the plunging region and reduce the importance of the ISCO, simply as a consequence of the MHD properties of the accreting fluid. Numerical simulations confirmed that the ISCO plays no important role in flow dynamics (see Figure 4, Figure 3 in Rule et al. 2025, and Lančová et al. 2025). The disk inner edge, as the location up to where the disk radiates as a blackbody, should be the (magneto)sonic point, rather than the ISCO (Sądowski, 2009). For low mass accretion rates, these coincide (Penna et al., 2012), but deviate strongly with increasing mass accretion rate. Results of simulations showed that the photosphere location and the density scale height may also differ significantly, in contrast to the assumption of analytical disk models, and even for low luminosities (e.g., Curd and Narayan, 2023; Liska et al., 2022; Mishra et al., 2022)
@@
The non-zero stress at ISCO may be parametrised by $\delta_{\rm j}$, defined as the ratio of the angular momentum passed back to the disk at ISCO to the angular momentum at ISCO, which approximately equals $\alpha(H/R)$ (Paczyński, 2000; Mummery and Balbus, 2023). However, there is a tension between what value it may reach (ranging from a fraction of a per cent to several per cent) and how much it may affect the spectra produced within the plunging region (compare results of Mummery et al. 2025 to those of Zhu et al. 2012). Numerical simulations yield different values, which may be caused by the differences in the simulated regime or the physics included. Those of Zhu et al. (2012) can be fitted with $\delta_{\rm j}\sim 0.5\%$, while Rule et al. (2025) report value of $\delta_{\rm j}\sim 5\%$, and the puffy disk simulations (which includes radiation and target higher mass accretion rates) give $\delta_{\rm j}\sim 16\%$ (Lančová et al., 2025). Furthermore, Lasota and Abramowicz (2024) argued that this value cannot be taken as a free parameter, since the solution would lead to an unphysical velocity on the BH horizon.
@@
Using an extended trans-ISCO thin disk model, Mummery et al. (2025) showed that the emission from the plunging region of a Schwarzschild BH mimics the effect of a high spin. This was also demonstrated by detailed spectral fitting of the X-ray spectra of the HMXB M33 X-7 with fullkerr, new model developed by Mummery et al. (2024) [^4] https://github.com/andymummeryastro/fullkerr, assuming $a_{*}=0$, obtaining $\delta_{\rm J}=0.078$. This fit was slightly better than that using kerrbb (i.e., at $\delta_{\rm J}=0$), which yielded $a_{*}=0.84\pm 0.05$. We note that Mummery et al. (2025) modelled the emission from $R>R_{\rm ISCO}$ by the standard model (Novikov and Thorne, 1973) except for the effect of the finite stress at the ISCO taken into account following Agol and Krolik (2000), but the issue of the thin disk model instabilities in a radiation-pressure-dominated regime is not discussed within their work.
@@
An associated effect is the capture of photons emitted within the ISCO. This modifies the standard limit of $a_{*{\rm max}}=0.998$ of Thorne (1974) to $a_{*{\rm max}}\lesssim 0.99$ for typical values of the stress at the ISCO found in numerical simulations (Mummery, 2025).
@@
An important question in the above models is whether they satisfy the $L\propto T_{\rm max}^{4}$ correlation found in many sources in the soft state (Gierliński and Done, 2004), where $T_{\rm max}$ is the maximum fitted disk temperature. This correlation suggests a constancy of the disk inner radius with changing luminosity, pointing to the importance of the ISCO. This would then be an important test of the validity of any modified disk model.
@@
On the other hand, the inner disk can remain aligned with the BH spin even in the presence of a BH spin-binary axis misalignment (Bardeen and Petterson, 1975). Such geometry, which is expected in case of a spin-orbit misalignment in X-ray binaries (Nixon et al., 2012; Marcel and Neilsen, 2021), appears to be supported in several cases by measurements of the X-ray polarization. In all cases in which there are both such measurements and resolved radio jets, the polarization angle is aligned with the jet position angle (e.g., Cyg X-1, Krawczynski et al. 2022; MAXI J1727.8–1613, Wood et al. 2024) or perpendicular (Cyg X-3, Veledina et al. 2024a, b). Assuming that jets are produced by the Blandford and Znajek (1977) mechanism (see Section 3.3), these results are compatible with the inner disk normal being aligned with the BH spin direction.
@@ 3.2 QPOs and spins
Yet another method derives the spin from QPOs. QPOs are narrow features observed in the X-ray power spectra of accreting compact objects. They are thought to reflect coherent, albeit transient, processes in the inner accretion flow. QPOs are typically categorized by frequency into low-frequency (LFQPOs; $\lesssim 30$ Hz) and high-frequency (HFQPOs; $\gtrsim 60$ Hz) types. Multiple models have been proposed to explain the most commonly observed ones at low-frequencies, ranging from geometric precession of the inner flow to disk oscillation modes and relativistic test-particle dynamics, see Ingram and Motta (2019) for a recent review. Among these, the Relativistic Precession Model (RPM) provides a framework for connecting all observed QPOs (both high and low-frequency) with the fundamental frequencies of particle motion around a spinning BH, offering a potential method to infer the BH’s spin and mass.
Current versions of this method are based on several theoretical models of disk oscillations occurring at a special radius of the disk and rely on pure test-particle dynamics. The existing models identify high-frequency QPOs with either a resonance between the orbital and radial epicyclic frequencies (Abramowicz and Kluźniak, 2001) or some other local disk oscillations. In particular, the QPO frequencies are matched by the RPM (Stella and Vietri, 1998, 1999; Stella et al., 1999). It offers a conceptually straightforward method to estimate BH spin by associating observed QPOs with fundamental frequencies of test-particle motion in the accretion flow: the periastron precession ($\nu_{\rm per}$), the orbital frequency ($\nu_{\phi}$), and the Lense-Thirring (Lense and Thirring, 1918) frequency, $\nu_{\rm LT}$, sometimes called the nodal frequency, $\nu_{\rm nod}$. See Motta et al. (2014) and Ingram and Motta (2019) for detailed reviews. It is crucial to distinguish the RPM model from the solid-body Lense-Thirring precession models that consider the global precession of the inner flow to explain the low-frequency QPOs (Fragile et al. 2007; Ingram et al. 2009; see discussion in Motta et al. 2018). Unlike solid-body precession models, the RPM does not rely on spin-orbit misalignment but instead on small perturbations to an otherwise circular, aligned disk. The key limitation of this model is that any perturbation of a small disk element (“test particle”) would be rapidly damped by interactions with the surrounding material, preventing the establishment of coherent relativistic precession. The RPM frequencies depend solely on the BH mass, spin, and the radius at which the oscillation originates – typically assumed to be the transition/truncation radius between the cold accretion disk and the hot accretion flow. In the innermost regions of a stellar-mass BH accretion flow, these frequencies lie in the $\sim 0.1$$10^{3}$ Hz range.
@@
While this method is seducing both by its simplicity and its consistency with the spins obtained from GW signals, we must remember that there is not yet a physical justification for the process producing the associated QPOs (unless one lies close to the ISCO, where, however, the dissipation is close to null). Furthermore, it requires the presence of slightly elliptical orbits (so that $\nu_{\rm per}>0$), which remains to be demonstrated, and non-zero BH spins (so that $\nu_{\rm LT}>0$). Moreover, in the case of neutron stars, where high-frequency QPOs are more common, the relationship between test-particle frequencies and the QPOs remains unclear (see, e.g., Méndez and Belloni 2007; van Doesburgh and van der Klis 2017).

### v3

@@ 3.1.2 The continuum method: standard models
Notes: A version of bhspec covering $a_{*}$ of 1-1–0.99 is available from Shane Davis upon request. The normalization should be kept at unity in all of the models except in bhspec, where it is tied to the source distance. Note that kerrbb is agnostic of the value of $L_{\rm d}/L_{\rm Edd}$ implied by its parameters, and thus the condition $L_{\rm d}/L_{\rm Edd}\lesssim 0.3$ needs to be checked a posteriori. The $f_{\rm col}$ of kerrbb2 is obtained by fitting its values in kerrbb to the corresponding spectra generated from bhspec.
A significant issue regarding the continuum method is the local rest-frame spectra of the disk. Table 1 lists the methods used by the main available models and the ranges of the main parameters, $a_{*}$, the viscosity parameter, $\alpha$, and the Eddington ratio, $L_{\rm d}/L_{\rm Edd}$, where $L_{\rm d}$ is the disk luminosity, and $L_{\rm Edd}$ is the Eddington luminosity for hydrogen. A commonly used model kerrbb (Li et al., 2005) approximates the local spectra as blackbodies with a color correction, $f_{\rm col}$ (Shimura and Takahara, 1995; Davis and El-Abd, 2019), where this factor is defined as the ratio of the color temperature to the effective one, $f_{\rm col}\equiv T_{\rm color}/T_{\rm eff}$. The value of $f_{\rm col}$ for a given disk is uncertain, though usually in the range of 1.5–2 (Davis and El-Abd, 2019). The same parametrization is used by the model kynbb [^1] https://projects.asu.cas.cz/stronggravity/kyn (Dovčiak et al., 2004b), which also allows the inner disk radius to be different from the ISCO radius, $R_{\rm ISCO}$.
@@
Finally, we notice that the simple nonrelativistic model of the disk continuum, diskbb (Mitsuda et al., 1984), is often used to roughly estimate the spin. The model gives an approximate (and proportional to the distance) estimate of the disk’s inner radius, which can be converted to the spin if the mass and distance are known. This model follows the formulation of Shakura and Sunyaev (1973) except that it does not include the zero-stress boundary condition at the ISCO, $(1-\sqrt{R_{\rm ISCO}/R})$. The distance, $D$, inclination, $i$, and the inner radius are related to the diskbb normalization, $N_{\rm dbb}$, by
@@ 3.1.3 The continuum method: problems and alternatives
An example of this class of models is the so-called magnetically elevated disks (Begelman and Pringle, 2007; Begelman and Silk, 2017; Mishra et al., 2020). Importantly, they have lower gas densities at the effective photosphere, which enhances electron scattering and could lead to a harder spectrum than from a disk without magnetic pressure support. For this reason, their average color correction could be significantly larger than the standard range of $f_{\rm col}\approx 1.5$22. No radiative transfer calculations of the spectra from magnetically elevated disks have, to our knowledge, been done yet. Still, there have been recent advancements in numerical simulations of accretion flows, including more realistic modeling of radiation transfer (e.g., Sądowski and Narayan, 2016; Huang et al., 2023; Fragile et al., 2023) as well as more sophisticated tools for generating spectral predictions (e.g., Kinch et al., 2019; Mills et al., 2024; Roth et al., 2025).
Alternatively, a significant part of the total pressure support could be provided by a radial magnetic field, in addition to the toroidal one (Lančová et al., 2019), resulting in a thicker disk, with a photosphere that is hotter than that of a standard disk, see Figure 2 in Wielgus et al. (2022). Radiative general relativistic magnetohydrodynamics (GRMHD) simulations of these so-called puffy disk models have been considered in the regime $\dot{M}_{\rm accr}\approx\dot{M}_{\rm Edd}$ (where we define $\dot{M}_{\rm Edd}\equiv L_{\rm Edd}/0.1c^{2}$). An example of the vertical structure of a puffy disk is shown in Figure 4, where the photosphere is located at around $z\approx r$ (solid white line at $\tau_{\rm T}=1$), while the disk scale height is $z\approx 0.1\,r$ (dashed white line). The presence of a hotter and geometrically thick photosphere reaching to the BH event horizon results in a significant overestimate of the BH spin. When the kerrbb and slimbh models were used on synthetic spectra obtained from the simulations assuming $a_{*}=0$, the fitted spin values ranged from $a_{*}\approx 0.5$ to $\approx$0.9 (Wielgus et al., 2022; Lančová et al., 2023). While these numerical simulations, like all others, come with numerous caveats, this case study is illuminating because it highlights how the more complex physical behavior seen in simulations (and potentially occurring in nature) can lead to significant biases in BH spin estimates, even when interpreted with state-of-the-art spectral models.
@@
Typically, the fitted coronal temperature is approximately the same as the inner disk temperature. This prevents the formation of a power-law spectral component due to Comptonization of the disk photons by those electrons. Instead, the overall spectral shape is modified. Together with the Galactic absorption being strong at $\lesssim$ 1 keV, these effects complicate placing effective spectral constraints on the existence of warm coronae in XRBs. Nevertheless, its presence was hinted at in some BH X-ray binaries (BH XRBs), particularly in the soft state of GRS 1915+105 (a BH X-ray binary with a low-mass donor, hereafter LMXB), which was fitted by a hybrid Comptonization model including a low-temperature, high optical depth thermal component (Zdziarski et al., 2001). A similar result was found for the very high state of GRO J1655–40 (Kubota et al., 2001).
Following those results, models with a warm coronae were fitted by Belczyński et al. (2024) and Zdziarski et al. (2024a, b, 2025) to the soft states of LMC X-1, Cyg X-1, M33 X-7, and GX 339–4. As we discuss in Section 4.1 below, this model provides statistically good fits and significantly alters the fitted values of the spin. However, the one-zone Comptonization model used so far for the warm corona (Zdziarski et al., 2020) does not account for the detailed corona structure presented in Gronkiewicz and Różańska (2020).
@@
Then, both analytical work and numerical simulations show that the viscous $\alpha$ parameter cannot be taken as constant, typically rising steeply near the ISCO and reaching a sharp maximum within the plunging region (Krolik, 1999; Agol and Krolik, 2000; Sorathia et al., 2012; Penna et al., 2013; Abramowicz et al., 1996; Jiang et al., 2019). Thus, this increase of stress will also affect the flow within the plunging region and reduce the importance of the ISCO, simply as a consequence of the MHD properties of the accreting fluid. Numerical simulations confirmed that the ISCO plays no important role in flow dynamics (see Figure 4, Figure 3 in Rule et al. 2025, and Lančová et al. 2025). The disk inner edge, as the location up to where the disk radiates as a blackbody, should be the (magneto)sonic point, rather than the ISCO (Sądowski, 2009). For low mass accretion rates, these coincide (Armitage et al., 2001; Penna et al., 2012), but deviate strongly with increasing mass accretion rate. Results of simulations showed that the photosphere location and the density scale height may also differ significantly, in contrast to the assumption of analytical disk models, and even for low luminosities (e.g., Curd and Narayan, 2023; Liska et al., 2022; Mishra et al., 2022).
@@
The non-zero stress at ISCO may be parametrized by $\delta_{\rm j}$, defined as the ratio of the angular momentum passed back to the disk at ISCO to the angular momentum at ISCO, which approximately equals $\alpha(H/R)$ (Paczyński, 2000; Mummery and Balbus, 2023). However, there is a tension between what value it may reach (ranging from a fraction of a percent to several percent) and how much it may affect the spectra produced within the plunging region (compare results of Mummery et al. 2025 to those of Zhu et al. 2012). Numerical simulations yield different values, which may be caused by the differences in the simulated regime or the physics included. Those of Zhu et al. (2012) can be fitted with $\delta_{\rm j}\sim 0.5\%$, while Rule et al. (2025) report value of $\delta_{\rm j}\sim 5\%$, and the puffy disk simulations (which includes radiation and target higher mass accretion rates) give $\delta_{\rm j}\sim 16\%$ (Lančová et al., 2025). Furthermore, Lasota and Abramowicz (2024) argued that this value cannot be taken as a free parameter, since the solution would lead to an unphysical velocity on the BH horizon.
@@
Using an extended trans-ISCO thin disk model, Mummery et al. (2025) showed that the emission from the plunging region of a Schwarzschild BH mimics the effect of a high spin. This was also demonstrated by detailed spectral fitting of the X-ray spectra of the HMXB M33 X-7 with fullkerr, a new model developed by Mummery et al. (2024) [^4] https://github.com/andymummeryastro/fullkerr, assuming $a_{*}=0$, obtaining $\delta_{\rm J}=0.078$. This fit was slightly better than that using kerrbb (i.e., at $\delta_{\rm J}=0$), which yielded $a_{*}=0.84\pm 0.05$. We note that Mummery et al. (2025) modeled the emission from $R>R_{\rm ISCO}$ by the standard model (Novikov and Thorne, 1973) except for the effect of the finite stress at the ISCO taken into account following Agol and Krolik (2000), but the issue of the thin disk model instabilities in a radiation-pressure-dominated regime is not discussed within their work.
@@
An associated effect is the presence of a limit on spin due to accretion. When magnetic fields are present, the stress continues to the horizon. This results in a strongly spin-dependent outward flux of angular momentum conveyed electromagnetically. Gammie et al. (2004) and Krolik et al. (2005) found that this effect limits the spin to $a_{*}\lesssim 0.9$. A weaker effect is the capture of photons emitted within the ISCO. The presence of magnetic fields modifies the standard limit of $a_{*{\rm max}}=0.998$ of Thorne (1974) to $a_{*{\rm max}}\lesssim 0.99$ for typical values of the stress at the ISCO found in numerical simulations (Mummery, 2025).
@@
An important question in the above models is whether they satisfy the $L\propto T_{\rm max}^{4}$ correlation found in many sources in the soft state (Gierliński and Done, 2004), where $T_{\rm max}$ is the maximum fitted disk temperature. This correlation suggests a constancy of the disk’s inner radius with changing luminosity, pointing to the importance of the ISCO. This would then be an important test of the validity of any modified disk model.
@@
On the other hand, the inner disk can remain aligned with the BH spin even in the presence of a BH spin-binary axis misalignment (Bardeen and Petterson, 1975). Such geometry, which is expected in case of a spin-orbit misalignment in X-ray binaries (Nixon et al., 2012; Marcel and Neilsen, 2021), appears, in several cases, to be supported by measurements of the X-ray polarization. In all cases in which there are both such measurements and resolved radio jets, the polarization angle is aligned with the jet position angle (e.g., Cyg X-1, Krawczynski et al. 2022; MAXI J1727.8–1613, Wood et al. 2024) or perpendicular (Cyg X-3, Veledina et al. 2024a, b). Assuming that the jets are produced by the spin-extraction mechanism (Blandford and Znajek 1977; see Section 3.3), these results are compatible with the inner disk normal being aligned with the BH spin direction.
@@ 3.2 QPOs and spins
Yet another method derives the spin from QPOs. QPOs are narrow features observed in the X-ray power spectra of accreting compact objects. They are thought to reflect coherent, albeit transient, processes in the inner accretion flow. QPOs are typically categorized by frequency into low-frequency ($\lesssim 30$ Hz) and high-frequency ($\gtrsim 60$ Hz) types. Multiple models have been proposed to explain the most commonly observed ones at low frequencies, ranging from geometric precession of the inner flow to disk oscillation modes and relativistic test-particle dynamics; see Ingram and Motta (2019) for a review. Among these, the Relativistic Precession Model (RPM) provides a framework for connecting all observed QPOs (both high and low-frequency) with the fundamental frequencies of particle motion around a spinning BH, offering a potential method to infer the BH spin and mass.
Current versions of this method are based on several theoretical models of disk oscillations occurring at special radii of the disk and rely on pure test-particle dynamics. The existing models identify high-frequency QPOs with either a resonance between the orbital and radial epicyclic frequencies (Abramowicz and Kluźniak, 2001) or some other local disk oscillations. In particular, the QPO frequencies are identified with those of the disk oscillations by the RPM (Stella and Vietri, 1998, 1999; Stella et al., 1999). It offers a conceptually straightforward method to estimate BH spin by associating observed QPOs with fundamental frequencies of test-particle motion in the accretion flow: the periastron precession ($\nu_{\rm per}$), the orbital frequency ($\nu_{\phi}$), and the Lense-Thirring (Lense and Thirring, 1918) frequency, $\nu_{\rm LT}$, sometimes called the nodal frequency, $\nu_{\rm nod}$. See Motta et al. (2014) and Ingram and Motta (2019) for detailed reviews. It is crucial to distinguish the RPM model from the solid-body Lense-Thirring precession models that consider the global precession of the inner flow to explain the low-frequency QPOs (Fragile et al. 2007; Ingram et al. 2009; see discussion in Motta et al. 2018). Unlike solid-body precession models, the RPM does not rely on spin-orbit misalignment but instead on small perturbations to an otherwise circular, aligned disk. The key limitation of this model is that any perturbation of a small disk element (“test particle”) would be rapidly damped by interactions with the surrounding material, preventing the establishment of coherent relativistic precession. The RPM frequencies depend solely on the BH mass, spin, and the radius at which the oscillation originates – typically assumed to be the transition/truncation radius between the cold accretion disk and the hot accretion flow. In the innermost regions of a stellar-mass BH accretion flow, these frequencies lie in the $\sim 0.1$$10^{3}$ Hz range.
@@
While this method is appealing both by its simplicity and its consistency with the spins obtained from GW signals, we must remember that there is no yet physical justification for the process producing the associated QPOs (unless one lies close to the ISCO, where, however, dissipation is close to null). Furthermore, it requires the presence of slightly elliptical orbits (so that $\nu_{\rm per}>0$), which remains to be demonstrated, and non-zero BH spins (so that $\nu_{\rm LT}>0$). Moreover, in the case of neutron stars, where high-frequency QPOs are more common, the relationship between test-particle frequencies and the QPOs remains unclear (see, e.g., Méndez and Belloni 2007; van Doesburgh and van der Klis 2017).